# Open-Set Domain Adaptation Under Background Distribution Shift: Challenges and A Provably Efficient Solution

### v2

@@ 1 Introduction
Adapting Machine Learning models to shifts in the data distribution is pivotal to ensuring their robustness and safety in real world applications (quinonero2008dataset; pmlr-v139-koh21a) for fields like healthcare (finlayson2021clinician), autonomous driving (filos2020can; wong2020identifying) and more broadly in computer vision (bendale2016towards). Maintaining robustness to distribution shift in classification problems includes both identifying familiar objects under new conditions, while detecting the long tail of object categories that has not been observed in the past. In this work we study the problem of adaptation when these two types of distribution shift occur concurrently. Specifically, we address the emergence of novel categories (panareda2017open) alongside shifting distributions of known classes, a scenario referred to as Open-Set Domain Adaptation (OSDA).
@@
Let our training data be $S_{{\mathcal{S}}}$. The core task here is to detect the novel classes in target data $S_{{\mathcal{T}}}$ collected under some conditions different from those observed at training time (panareda2017open) while maintaining good performance over the existing classes. The emergence of a novel class in itself constitutes a shift between $S_{{\mathcal{S}}}$ and $S_{{\mathcal{T}}}$, however, in real world scenarios it is likely not the only distribution shift that occurs (garg22adaptation; wald2023birds). Our study concentrates on scenarios where the novel class exhibits a notable degree of separability from the known classes, and a distribution shift with overlapping supports between source and target distributions exists within the known classes, which we refer to as ‘background shift’. Figure 1 depicts a scenario of background shift within indoor scene images taken from the SUN397 dataset (jianxiong2010sun397) while the novel class belongs to an outdoor scene category. Arguably, this is a very common case in domains such as healthcare, e.g. classifying histopathology slides for known and novel tumor types, where the background shift is due to differences in geographic locations (pocevivciute2025out).
@@
Several aspects of our work are understudied in the OSDA literature; (i) formally characterizing the shifts of interest; (ii) developing methods with formal guarantees (finite sample bounds in our case) that are also effective in practice; (iii) examining the effect of the novel class size on performance. Most research primarily address open-set recognition instead of OSDA because they assume there is no distribution shift for known classes (ge2017openmax; Neal_2018_ECCV; Lin_2019; Chen_2021; zeng-etal-2021-modeling; vaze2022openset; Esmaeilpour_Liu_Robertson_Shu_2022). Other methods tackle OSDA without formally characterizing ‘domain shifts’ (liang2021reallyneedaccesssource; li2021domainconsensus). They typically validate their methods empirically on benchmarks such as Office-Home (VenkateswaraECP17deephashing) or VisDA (visda2017), where the shift is simulated by changing the domain from real to sketch images or from synthetic to real world images. However, since these methods lack formal guarantees and rely on limited types of simulated shifts, their effectiveness beyond these datasets remain uncertain. garg22adaptation tackle OSDA under shifting class probabilities in $S_{{\mathcal{T}}}$ (label shift) and show improved results compared to prior methods. To this end, our theoretical and empirical results emphasize the regimes where our method is expected to outperform alternatives that do not address background shift in OSDA properly. In particular, we show that our approach is more effective when the novel class constitutes a relatively small proportion of the target distribution—an important scenario in practice. Reliably identifying novel classes enables updating models or deferring to human oversight in safety-critical settings such as healthcare and autonomous driving (amodei2016concrete; hendrycks2016baseline).
@@
  • Leverage and improve constrained learning rule for OSDA: We develop CoLOR (Constrained Learning for Open-set Recognition), a scalable method to train models for OSDA (section˜5). Our work builds on a method developed for a special case of our setting, where small models are trained and there are no known classes (wald2023birds). Our architectural and algorithmic improvements enable us to work effectively with large models and complex, high-dimensional datasets such as images and text. Our theoretical contributions, explain why and when the method performs favorably in the large scale regime.
  • Formal results on identifiability under background shift: We characterize sufficient conditions for identifiability of the novel class and analyze sample efficiency of CoLOR on a problem that captures key aspects of OSDA with expressive models (section˜4.2). While prior work (garg22adaptation) shows that OSDA under label shift among non-novel instances is solvable under weaker assumptions than our separability condition, we prove that these assumptions fail to handle the more general case of background shift (see Lemma 1). Using linear–Gaussian framework, we characterize problem regimes where CoLOR significantly outperforms a standard domain-discriminator baseline, especially when novel classes are rare.
  • Comprehensive empirical validation across diverse data modalities and datasets: Extensive experiments across image and text classification (section˜6) confirm our theoretical findings. We observe that background shift substantially degrades OSDA baselines, whereas our enhanced adaptation of the constrained learning rule, CoLOR, achieves superior performance on CIFAR100 (Krizhevsky09learningmultiplecifar), Amazon Reviews (ni_justifying_2019) and SUN397 (jianxiong2010sun397). Notably, CoLOR remains robust even when the novel class constitutes a small portion of $S_{{\mathcal{T}}}$, where other methods tend to struggle.
@@ 2 Related Work
Out-of-Distribution (OOD) Detection and Open Set Recognition: Most works on detecting novel instances aim to locate novelties within $S_{{\mathcal{T}}}$ without adapting to it. To this end, methods use $S_{{\mathcal{S}}}$ alone to determine an “anomaly score” $s({\mathbf{x}})$ (i.e. a score that ranks examples as more or less likely to be novel) (ruff2021unifying; Esmaeilpour_Liu_Robertson_Shu_2022). Related to this setting is Open Set Recognition (OSR) (bendale2016towards; ge2017openmax; liu2018open; xu2019open; saito2018open; vaze2022openset), where the goal is again to detect the novelties (while classifying known classes correctly), but the learner is given access to $S_{{\mathcal{T}}}$ without the labels. This essentially grants us the ability to adapt $s({\mathbf{x}})$ to the specific novel class presented in $S_{{\mathcal{T}}}$. Some OSR approaches use generative models to generate synthetic images using known instances to further improve $s({\mathbf{x}})$ (Neal_2018_ECCV; Kong_2021_ICCV_opengan). However, both OSR and OOD detection methods suffer from degraded performance when there is a distribution shift between $P_{{\mathcal{S}}}$ and the distribution of known classes under $P_{{\mathcal{T}}}$ ($P_{{\mathcal{T}},[k]}$) (scholkopf2001estimating; cao2023anomalydetectiondistributionshift), hence methods that are adaptable to such shifts have been developed.
@@
Open Set Domain Adaptation: Works that seek to adapt to $P_{{\mathcal{T}},[k]}$, while correctly classifying known instances and detecting novel ones, fall into this category. Recent works on OSDA primarily focus on specific domain shifts from real images to sketches or real to synthetic (wen24crossdomain; panareda2017open; choe2024osdasemanticsegmentation; Hur2023unida; Zhu2023unida). However, the narrow focus on these shifts limits their generalizability to broader shifts, such as background shift. garg22adaptation show that these methods may not be versatile across datasets and propose a sophisticated OSDA method specifically to handle label shift. Additionally, prior works have leveraged multi-modal foundation models like ChatGPT, DALL-E and CLIP for OSDA (qu2023lmc; Esmaeilpour_Liu_Robertson_Shu_2022). However, since their pretraining data is not publicly accessible, it is infeasible to curate truly novel classes or distribution shifts. This makes it difficult to ensure fair evaluation of these methods on open-set problems. Most OSDA methods lack formal guarantees or a clear characterization of the shifts they address. In this context, well-studied Positive and Unlabeled (PU) learning problems offer a useful framework to develop more generalizable methods with guarantees (garg22adaptation; wald2023birds).
@@
PU-Learning Under Distribution Shift: It is formally equivalent to OSDA where k=1. In the absence of distribution shift, effective solutions with formal guarantees already exist, primarily based on adjustments to “Domain Discriminator” models (that are trained to distinguish source and target distributions i.e. $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}}}$ respectively) (garg2021mixture; duplessis2014analysis; blanchard2010semi; elkan2008learning). Under distribution shift, domain discriminator type methods still have guarantees when given infinite data (gerych2022recovering). In practice, however, we will see that they underperform (footnote˜11), even when the separability assumption (see assumption˜1) holds. garg22adaptation propose a $k$-way PU learning based method for identifying unseen classes under special case of label shift among $k$ known classes, given certain assumptions. We show that these assumptions are insufficient to handle the more general case of background shift, where additional conditions are required for reliable recovery. Furthermore, wald2023birds propose an algorithm with finite-sample guarantees and demonstrate its performance on small-scale models. In contrast, our formal analysis targets the large-scale, overparameterized learning regime, where their results no longer apply (belkin2019reconciling). We extend these ideas to the setting of OSDA under background shift.
@@ 3.1 Problem Definition
The success of an OSDA model may be determined by its classification accuracy on data from $P_{{\mathcal{T}}}$, $\mathcal{R}^{l_{01}}_{\mathcal{T}}(h)=\mathbb{E}_{y,{\mathbf{x}}\sim P_{{\mathcal{T}}}}[\mathrm{arg}\max_{\hat{y}}\{h({\mathbf{x}})_{\hat{y}}\}\neq y]$, but since we often wish to emphasize novelty detection and not treat the novel class like any other known class, several other notions of success have been defined in the literature. Our theory focuses on the simple case in which $k=1$, where such considerations are unnecessary, and standard metrics for binary classification can be used. In our experiments, as we mention again in section˜6, we also use the OSCR metric of dhamija2018oscr. Below we formally define the problem setting and other performance metrics we use for our theory.
@@ 3.2 Necessary and Sufficient Assumptions for OSDA under Background Shift
Without any assumptions on the relationship between $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}}}$ it is impossible to obtain a guarantee of better than chance accuracy. This holds even under strong simplifying assumptions, such as the existence of a model that solves the problem without errors: that is, $\mathcal{R}^{l_{01}}_{\mathcal{T}}(h^{*})=0$ for some hypothesis $h^{*}$ 222see Prop. 1 in garg22adaptation, Prop. 2.1 in wald2023birds, or discussion in bekker2019beyond for the results of this flavor..
@@
Separability and background shift characteristics in known classes. The separability assumption is required only for the novel class $Y=k+1$ vs. the known ones, we do not explicitly limit the $k$ known classes to be separable amongst themselves. Such an assumption would have placed the shift between $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}},[k]}$ purely in the realm of covariate shift (shimodaira2000improving). However, background shift with assumption˜1 facilitates more general forms of label shift and covariate shift. In our experiments we mostly create distribution shifts such that $P_{{\mathcal{T}},[k]}(X|Y)\neq P_{{\mathcal{S}}}(X~|~Y)$ which follow the definition of background shift that are underexplored in existing works.
Insufficiency of less stringent assumptions. While generalization bounds for domain adaptation are well known from seminal works such as bendavid2010adaptation, they are less common in the Open Set learning literature, hence let us focus on this aspect of our problem, i.e. detecting the novel class. To the best of our knowledge the only characterized sufficient and necessary conditions for (non-separable) OSDA are those in garg22adaptation, which hold only for label shift, where it is assumed that $P_{{\mathcal{S}}}(X~|~Y=y)=P_{{\mathcal{T}}}(X~|~Y=y)$ for all $y\in{[k]}$. For this special case, they propose two assumptions which are sufficient (when added on top of the label-shift assumption) to guarantee $h^{*}$ can be learned from observed data. Their first assumption is (Strong Positivity): there exists $X_{sep}\in{{\mathcal{X}}}$ such that $P_{{\mathcal{T}},k+1}(X_{sep})=0$ and the matrix $P_{{\mathcal{S}}}({\mathbf{x}}~|~y)]_{{\mathbf{x}}\in{X_{sep}},y\in{[k]}}$ is full rank and diagonal. We show that once more general shifts than label shifts are allowed, e.g. background/covariate shift, this condition is no longer sufficient and in fact no algorithm can guarantee better-than-chance detection. The proof for this is in A.3.
@@
The second assumption proposed in garg22adaptation is separability as defined in assumption˜1; however, since their assumption is combined with the label shift assumption, the methods they develop are tailored to that scenario and do not apply to our problem.  [^3] Lemma 1 has a similar flavor to Prop. 3.1 in wald2023birds, but it is stronger. Our result shows impossibility under an additional assumption of strong positivity. This is significant in the context we consider here, as garg22adaptation give guarantees on OSDA with label shift under this strong positivity assumption, but our lemma shows that this is impossible under the background shifts we consider.
Assumption˜1 guarantees that given an infinitely large sample and an optimization oracle, we can learn to identify the novelties. It is also desirable to have guarantees on the required sample size. In this paper we focus on the finite sample guarantees of detecting the novel class in the overparameterized regime, since results on classification of remaining $k$ non-novel classes can be obtained from prior work on domain adaptation, see for instance ben2010theory and the survey of redko2020survey. This is to formally study and understand best practices in learning expressive models for open-set problems under background shift.

### v3

@@ 1 Introduction
Adapting Machine Learning models to shifts in the data distribution is pivotal to ensuring their robustness and safety in real world applications (Quinonero-Candela et al., 2008; Koh et al., 2021) for fields like healthcare (Finlayson et al., 2021), autonomous driving (Filos et al., 2020; Wong et al., 2020) and more broadly in computer vision (Bendale and Boult, 2016). Maintaining robustness to distribution shift in classification problems includes both identifying familiar objects under new conditions, while detecting the long tail of object categories that has not been observed in the past. In this work we study the problem of adaptation when these two types of distribution shift occur concurrently. Specifically, we address the emergence of novel categories (Panareda Busto and Gall, 2017) alongside shifting distributions of known classes, a scenario referred to as Open-Set Domain Adaptation (OSDA).
@@
Let our training data be $S_{{\mathcal{S}}}$. The core task here is to detect the novel classes in target data $S_{{\mathcal{T}}}$ collected under some conditions different from those observed at training time (Panareda Busto and Gall, 2017) while maintaining good performance over the existing classes. The emergence of a novel class in itself constitutes a shift between $S_{{\mathcal{S}}}$ and $S_{{\mathcal{T}}}$, however, in real world scenarios it is likely not the only distribution shift that occurs (Garg et al., 2022; Wald and Saria, 2023). Our study concentrates on scenarios where the novel class exhibits a notable degree of separability from the known classes, and a distribution shift with overlapping supports between source and target distributions exists within the known classes, which we refer to as ‘background shift’. Figure 1 depicts a scenario of background shift within indoor scene images taken from the SUN397 dataset (Xiao et al., 2010) while the novel class belongs to an outdoor scene category. Arguably, this is a very common case in domains such as healthcare, e.g. classifying histopathology slides for known and novel tumor types, where the background shift is due to differences in geographic locations (Pocevičiūtė et al., 2025).
@@
Several aspects of our work are understudied in the OSDA literature; (i) formally characterizing the shifts of interest; (ii) developing methods with formal guarantees (finite sample bounds in our case) that are also effective in practice; (iii) examining the effect of the novel class size on performance. Most research primarily address open-set recognition instead of OSDA because they assume there is no distribution shift for known classes (Ge et al., 2017; Neal et al., 2018; Lin and Xu, 2019; Chen et al., 2021; Zeng et al., 2021; Vaze et al., 2022; Esmaeilpour et al., 2022). Other methods tackle OSDA without formally characterizing ‘domain shifts’ (Liang et al., 2021; Li et al., 2021). They typically validate their methods empirically on benchmarks such as Office-Home (Venkateswara et al., 2017) or VisDA (Peng et al., 2017), where the shift is simulated by changing the domain from real to sketch images or from synthetic to real world images. However, since these methods lack formal guarantees and rely on limited types of simulated shifts, their effectiveness beyond these datasets remain uncertain. Garg et al. (2022) tackle OSDA under shifting class probabilities in $S_{{\mathcal{T}}}$ (label shift) and show improved results compared to prior methods. To this end, our theoretical and empirical results emphasize the regimes where our method is expected to outperform alternatives that do not address background shift in OSDA properly. In particular, we show that our approach is more effective when the novel class constitutes a relatively small proportion of the target distribution—an important scenario in practice. Reliably identifying novel classes enables updating models or deferring to human oversight in safety-critical settings such as healthcare and autonomous driving (Amodei et al., 2016; Hendrycks and Gimpel, 2016).
@@
  • Leverage and improve constrained learning rule for OSDA: We develop CoLOR (Constrained Learning for Open-set Recognition), a scalable method to train models for OSDA (section˜5). Our work builds on a method developed for a special case of our setting, where small models are trained and there are no known classes (Wald and Saria, 2023). Our architectural and algorithmic improvements enable us to work effectively with large models and complex, high-dimensional datasets such as images and text. Our theoretical contributions, explain why and when the method performs favorably in the large scale regime.
  • Formal results on identifiability under background shift: We characterize sufficient conditions for identifiability of the novel class and analyze sample efficiency of CoLOR on a problem that captures key aspects of OSDA with expressive models (section˜4.2). While prior work (Garg et al., 2022) shows that OSDA under label shift among non-novel instances is solvable under weaker assumptions than our separability condition, we prove that these assumptions fail to handle the more general case of background shift (see Lemma 1). Using linear–Gaussian framework, we characterize problem regimes where CoLOR significantly outperforms a standard domain-discriminator baseline, especially when novel classes are rare.
  • Comprehensive empirical validation across diverse data modalities and datasets: Extensive experiments across image and text classification (section˜6) confirm our theoretical findings. We observe that background shift substantially degrades OSDA baselines, whereas our enhanced adaptation of the constrained learning rule, CoLOR, achieves superior performance on CIFAR100 (Krizhevsky, 2009), Amazon Reviews (Ni et al., 2019) and SUN397 (Xiao et al., 2010). Notably, CoLOR remains robust even when the novel class constitutes a small portion of $S_{{\mathcal{T}}}$, where other methods tend to struggle.
@@ 2 Related Work
Out-of-Distribution (OOD) Detection and Open Set Recognition: Most works on detecting novel instances aim to locate novelties within $S_{{\mathcal{T}}}$ without adapting to it. To this end, methods use $S_{{\mathcal{S}}}$ alone to determine an “anomaly score” $s({\mathbf{x}})$ (i.e. a score that ranks examples as more or less likely to be novel) (Ruff et al., 2021; Esmaeilpour et al., 2022). Related to this setting is Open Set Recognition (OSR) (Bendale and Boult, 2016; Ge et al., 2017; Liu et al., 2018; Xu et al., 2019; Saito et al., 2018; Vaze et al., 2022), where the goal is again to detect the novelties (while classifying known classes correctly), but the learner is given access to $S_{{\mathcal{T}}}$ without the labels. This essentially grants us the ability to adapt $s({\mathbf{x}})$ to the specific novel class presented in $S_{{\mathcal{T}}}$. Some OSR approaches use generative models to generate synthetic images using known instances to further improve $s({\mathbf{x}})$ (Neal et al., 2018; Kong and Ramanan, 2021). However, both OSR and OOD detection methods suffer from degraded performance when there is a distribution shift between $P_{{\mathcal{S}}}$ and the distribution of known classes under $P_{{\mathcal{T}}}$ ($P_{{\mathcal{T}},[k]}$) (Schölkopf et al., 2001; Cao et al., 2023), hence methods that are adaptable to such shifts have been developed.
@@
Open Set Domain Adaptation: Works that seek to adapt to $P_{{\mathcal{T}},[k]}$, while correctly classifying known instances and detecting novel ones, fall into this category. Recent works on OSDA primarily focus on specific domain shifts from real images to sketches or real to synthetic (Wen and Brbic, 2024; Panareda Busto and Gall, 2017; Choe et al., 2024; Hur et al., 2023; Zhu et al., 2023). However, the narrow focus on these shifts limits their generalizability to broader shifts, such as background shift. Garg et al. (2022) show that these methods may not be versatile across datasets and propose a sophisticated OSDA method specifically to handle label shift. Additionally, prior works have leveraged multi-modal foundation models like ChatGPT, DALL-E and CLIP for OSDA (Qu et al., 2023; Esmaeilpour et al., 2022). However, since their pretraining data is not publicly accessible, it is infeasible to curate truly novel classes or distribution shifts. This makes it difficult to ensure fair evaluation of these methods on open-set problems. Most OSDA methods lack formal guarantees or a clear characterization of the shifts they address. In this context, well-studied Positive and Unlabeled (PU) learning problems offer a useful framework to develop more generalizable methods with guarantees (Garg et al., 2022; Wald and Saria, 2023).
@@
PU-Learning Under Distribution Shift: It is formally equivalent to OSDA where k=1. In the absence of distribution shift, effective solutions with formal guarantees already exist, primarily based on adjustments to “Domain Discriminator” models (that are trained to distinguish source and target distributions i.e. $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}}}$ respectively) (Garg et al., 2021; du Plessis et al., 2014; Blanchard et al., 2010; Elkan and Noto, 2008). Under distribution shift, domain discriminator type methods still have guarantees when given infinite data (Gerych et al., 2022). In practice, however, we will see that they underperform (footnote˜11), even when the separability assumption (see assumption˜1) holds. Garg et al. (2022) propose a $k$-way PU learning based method for identifying unseen classes under special case of label shift among $k$ known classes, given certain assumptions. We show that these assumptions are insufficient to handle the more general case of background shift, where additional conditions are required for reliable recovery. Furthermore, Wald and Saria (2023) propose an algorithm with finite-sample guarantees and demonstrate its performance on small-scale models. In contrast, our formal analysis targets the large-scale, overparameterized learning regime, where their results no longer apply (Belkin et al., 2019). We extend these ideas to the setting of OSDA under background shift.
@@ 3.1 Problem Definition
The success of an OSDA model may be determined by its classification accuracy on data from $P_{{\mathcal{T}}}$, $\mathcal{R}^{l_{01}}_{\mathcal{T}}(h)=\mathbb{E}_{y,{\mathbf{x}}\sim P_{{\mathcal{T}}}}[\mathrm{arg}\max_{\hat{y}}\{h({\mathbf{x}})_{\hat{y}}\}\neq y]$, but since we often wish to emphasize novelty detection and not treat the novel class like any other known class, several other notions of success have been defined in the literature. Our theory focuses on the simple case in which $k=1$, where such considerations are unnecessary, and standard metrics for binary classification can be used. In our experiments, as we mention again in section˜6, we also use the OSCR metric of Dhamija et al. (2018). Below we formally define the problem setting and other performance metrics we use for our theory.
@@ 3.2 Necessary and Sufficient Assumptions for OSDA under Background Shift
Without any assumptions on the relationship between $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}}}$ it is impossible to obtain a guarantee of better than chance accuracy. This holds even under strong simplifying assumptions, such as the existence of a model that solves the problem without errors: that is, $\mathcal{R}^{l_{01}}_{\mathcal{T}}(h^{*})=0$ for some hypothesis $h^{*}$ 222see Prop. 1 in Garg et al. (2022), Prop. 2.1 in Wald and Saria (2023), or discussion in Bekker et al. (2019) for the results of this flavor..
@@
Separability and background shift characteristics in known classes. The separability assumption is required only for the novel class $Y=k+1$ vs. the known ones, we do not explicitly limit the $k$ known classes to be separable amongst themselves. Such an assumption would have placed the shift between $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}},[k]}$ purely in the realm of covariate shift (Shimodaira, 2000). However, background shift with assumption˜1 facilitates more general forms of label shift and covariate shift. In our experiments we mostly create distribution shifts such that $P_{{\mathcal{T}},[k]}(X|Y)\neq P_{{\mathcal{S}}}(X~|~Y)$ which follow the definition of background shift that are underexplored in existing works.
Insufficiency of less stringent assumptions. While generalization bounds for domain adaptation are well known from seminal works such as Ben-David et al. (2010b), they are less common in the Open Set learning literature, hence let us focus on this aspect of our problem, i.e. detecting the novel class. To the best of our knowledge the only characterized sufficient and necessary conditions for (non-separable) OSDA are those in Garg et al. (2022), which hold only for label shift, where it is assumed that $P_{{\mathcal{S}}}(X~|~Y=y)=P_{{\mathcal{T}}}(X~|~Y=y)$ for all $y\in{[k]}$. For this special case, they propose two assumptions which are sufficient (when added on top of the label-shift assumption) to guarantee $h^{*}$ can be learned from observed data. Their first assumption is (Strong Positivity): there exists $X_{sep}\in{{\mathcal{X}}}$ such that $P_{{\mathcal{T}},k+1}(X_{sep})=0$ and the matrix $P_{{\mathcal{S}}}({\mathbf{x}}~|~y)]_{{\mathbf{x}}\in{X_{sep}},y\in{[k]}}$ is full rank and diagonal. We show that once more general shifts than label shifts are allowed, e.g. background/covariate shift, this condition is no longer sufficient and in fact no algorithm can guarantee better-than-chance detection. The proof for this is in A.3.
@@
The second assumption proposed in Garg et al. (2022) is separability as defined in assumption˜1; however, since their assumption is combined with the label shift assumption, the methods they develop are tailored to that scenario and do not apply to our problem.  [^3] Lemma 1 has a similar flavor to Prop. 3.1 in Wald and Saria (2023), but it is stronger. Our result shows impossibility under an additional assumption of strong positivity. This is significant in the context we consider here, as Garg et al. (2022) give guarantees on OSDA with label shift under this strong positivity assumption, but our lemma shows that this is impossible under the background shifts we consider.
Assumption˜1 guarantees that given an infinitely large sample and an optimization oracle, we can learn to identify the novelties. It is also desirable to have guarantees on the required sample size. In this paper we focus on the finite sample guarantees of detecting the novel class in the overparameterized regime, since results on classification of remaining $k$ non-novel classes can be obtained from prior work on domain adaptation, see for instance Ben-David et al. (2010a) and the survey of Redko et al. (2020). This is to formally study and understand best practices in learning expressive models for open-set problems under background shift.